# Classical simulation of non-Gaussian bosonic circuits

### v1

@@ 4.2 Computing the overlap between Gaussian states
(a) The input to $\mathsf{overlap}$ consists of descriptions of two Gaussian states $\psi(\Delta_{1}),\psi(\Delta_{2})$ which includes $\alpha_{1},\alpha_{2}\in\mathbb{C}^{n}$ and the overlaps $r_{j}=\langle\alpha_{j},\psi(\Delta_{j})\rangle,j\in[2]$.
(b) The algorithm computes $\lambda\in\mathbb{C}^{n}$ and $\vartheta\in\mathbb{R}$ such that $D(\lambda)|\alpha_{1}\rangle=e^{i\vartheta}|\alpha_{2}\rangle$. This implies that $\langle\alpha_{2},D(\lambda)\alpha_{1}\rangle=e^{i\vartheta}$ and $\langle\psi(\Delta_{2}),D(\lambda)\alpha_{1}\rangle=e^{i\vartheta}\overline{r_{2}}$ are known.
(c) Lastly, the subroutine $\mathsf{overlaptriple}$ is applied to complete a triangle. This amounts to computing the inner product $w=\langle\psi(\Delta_{1}),\psi(\Delta_{2})\rangle$ which the algorithm returns.
Figure 4: An illustration of the algorithm $\mathsf{overlap}$.
@@ 4.3 Computing descriptions of evolved states
$\alpha^{\prime}_{k}$ $=\alpha_{k}\qquad\text{for}\qquad k\neq j\ .$ (100)
Because of this, we build an algorithmic subroutine called $\mathsf{squeezing}$ which, given the description of a Gaussian state $|\psi\rangle$, a parameter $z\in(0,+\infty]$ and $j\in[n]$, returns the description of the squeezed state $S_{j}(z)|\psi\rangle$. We call this subroutine in the algorithm $\mathsf{applyunitary}$ when $U$ is a single-mode squeezing operator.
First, we give the subroutine $\mathsf{squeezing}$, including pseudocode in Fig. 5 and associated claims in Lemma 4.6, as well as a pictorial description of the algorithm in Fig. 6
@@
1:$S\leftarrow S$ given in the ‘single-mode squeezing’ row of Table 1
2:$\operatorname{\Gamma}^{\prime}\leftarrow S\Gamma S^{T}$triangleright covariance matrix of $S_{j}(z)|\psi\rangle$
3:$\alpha^{\prime}\leftarrow\alpha$
4:$\alpha^{\prime}_{j}\leftarrow\alpha_{j}\cosh(z)-\overline{\alpha_{j}}\sinh(z)$
5:$d_{1},d_{2},d_{3}\leftarrow\hat{d}(\alpha^{\prime})$triangleright displacement vectors of $S_{j}(z)|\alpha\rangle,|\alpha^{\prime}\rangle,S_{j}(z)|\psi\rangle$
6:$(\Gamma_{1},\Gamma_{2},\Gamma_{3})\leftarrow(I,I,\Gamma^{\prime})$triangleright covariance matrices of $(S_{j}(z)|\alpha\rangle,|\alpha\rangle,S_{j}(z)|\psi\rangle)$
7:$u\leftarrow\overline{r}$triangleright compute $\langle S_{j}(z)\psi,S_{j}(z)\alpha\rangle$
8:$v\leftarrow 1/\sqrt{\cosh(z)}$triangleright compute $\langle S_{j}(z)\alpha,\alpha^{\prime}\rangle$
9:$r^{\prime}\leftarrow\mathsf{overlaptriple}(\Gamma_{1},\Gamma_{2},\Gamma_{3},d_{1},d_{2},d_{3},u,v,0)$triangleright compute $\langle\alpha^{\prime},S_{j}(z)\psi\rangle$
10:return $(\operatorname{\Gamma}^{\prime},\alpha^{\prime},r^{\prime})$
Figure 5: Given the description $\Delta\in\operatorname{\mathsf{Desc}}_{n}$ of a Gaussian state $|\psi\rangle\in\mathrm{Gauss}_{n}$, a parameter $z\in(0,+\infty)$ and $j\in[n]$, the algorithm $\mathsf{squeezing}$ returns the description of the state $S_{j}(z)|\psi(\Delta)\rangle$.
@@
The algorithm $\mathsf{squeezing}:\operatorname{\mathsf{Desc}}_{n}\times(0,+\infty)\times[n]\rightarrow\operatorname{\mathsf{Desc}}_{n}$ given in Fig. 5 runs in time $O(n^{3})$. The algorithm is such that
$|\psi(\mathsf{squeezing}(\operatorname{\Delta},z,j))\rangle=S_{j}(z)|\psi(\operatorname{\Delta})\rangle\qquad\text{for all}\qquad\operatorname{\Delta}\in\operatorname{\mathsf{Desc}}_{n},z\in\mathbb{R}\text{ and }j\in[n].$ (101)
That is, the output of the algorithm is a description of the state $S_{j}(z)|\psi(\Delta)\rangle$.
@@
(a) The input to the algorithm $\mathsf{squeezing}$ is the description of a Gaussian state $\psi$ which includes the overlap $r=\langle\alpha,\psi\rangle$, as well as $z\in\mathbb{R}$ and an index $j\in[n]$ associated to a single-mode squeezing operation $S_{j}(z)$.
(b) By unitarity of $S_{j}(z)$, the input data provides the inner product $\langle S_{j}(z)\alpha,S_{j}(z)\psi\rangle=\langle\alpha,\psi\rangle=r$.
(c) The reference state of the evolved state $S_{j}(z)|\psi\rangle$ is the coherent state $|\alpha^{\prime}\rangle$ with $\alpha^{\prime}\in\mathbb{C}^{n}$ such that $S(z)|\alpha\rangle=D(\alpha^{\prime})S(z)|0\rangle$. This is due to $S_{j}(z)|\psi\rangle$ having the same displacement vector $d(\alpha^{\prime})$ as $S_{j}(z)|\alpha\rangle$.
(d) The inner product $\langle\alpha^{\prime},S_{j}(z)\alpha\rangle$ is computed using Eq. 113.
(e) In the last step, the subroutine $\mathsf{overlaptriple}$ is used to compute $r^{\prime}=\langle\alpha^{\prime},S_{j}(z)\psi\rangle$. Thus, $(\operatorname{\Gamma}^{\prime},\alpha^{\prime},r^{\prime})$ is the description of $S_{j}(z)|\psi\rangle$.
Figure 6: An illustration of the algorithm $\mathsf{squeezing}$ laid out in Fig. 5. Dashed lines correspond to inner products between the states at the vertices which are known to be non-zero but have not been computed.
@@
where $\alpha^{\prime}$ is the vector given in Eqs. 99 and 100 and initialized in Lines 3 and 4. The state $S_{j}(z)|\alpha\rangle$ has displacement vector $\hat{d}(\alpha^{\prime})$. Since $|\alpha\rangle$ and $|\psi(\Delta)\rangle$ have the same displacement (by the definition of $\Delta$) and $S_{j}(z)$ maps states with the same displacement to states with the same displacement, then the state $S_{j}(z)|\psi(\Delta)\rangle$ has the same displacement $\hat{d}(\alpha^{\prime})$ as $|\alpha^{\prime}\rangle$. Then, $\Delta^{\prime}$ satisfies Item (i) in Definition 4.1.
It remains to show that the description $\Delta^{\prime}$ satisfies Item (ii) in Definition 4.1. This corresponds to showing that $r^{\prime}=\langle\alpha,S_{j}(z)\psi\rangle$. In Lines 5 and 6 we set
$d_{j}$ $=\hat{d}(\alpha^{\prime})\qquad\text{for}\qquad j\in[3]\ ,$ (106)
$(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ $=(I,I,\Gamma^{\prime})\ ,$ (107)
@@
which we set as $r^{\prime}$ in Line 9. Hence, the description $\Delta^{\prime}$ satisfies Item (ii) in Definition 4.1.
@@
We give pseudocode for the algorithm $\mathsf{applyunitary}$ in Fig. 7 and in Theorem 4.7 we prove that it outputs the description of a Gaussian state evolved with a Gaussian unitary operation. We also give a diagram of the algorithm in Fig. 8 when the operation is a displacement, phase shit or beamsplitter. When the operation is a single-mode squeezing operator, we refer to Fig. 6 for a diagram of the algorithm.
@@
2:     $(z,j)\leftarrow\Delta_{U}$triangleright $\Delta_{U}=(z,j)$ is a description of $U=S_{j}(z)$
@@
6:     $\beta\leftarrow\Delta_{U}$triangleright $\Delta_{U}=\beta$ is a description of $U=D(\beta)$
@@
11:     $(\phi,j)\leftarrow\Delta_{U}$triangleright $\Delta_{U}=(\phi,j)$ is a description of $U=F_{j}(\phi)$
@@
16:     $(\omega,j,k)\leftarrow\Delta_{U}$triangleright $\Delta_{U}=(\omega,j,k)$ is a description of $U=B_{j,k}(\omega)$
@@
20:$\operatorname{\Gamma}^{\prime}\leftarrow S\Gamma S^{T}$triangleright covariance matrix of $U|\psi\rangle$
@@
Figure 7: Given the description $\Delta\in\operatorname{\mathsf{Desc}}_{n}$ of a Gaussian state $\psi\in\mathrm{Gauss}_{n}$, a description $\Delta_{U}$ of a Gaussian unitary $U$ and a label $\ell\in\{\mathsf{displacement},\mathsf{phaseshift},\allowbreak\mathsf{beamsplitter},\mathsf{squeezing}\}$, the algorithm $\mathsf{applyunitary}$ returns the description of the state $U|\psi(\Delta)\rangle$.
@@
The algorithm $\mathsf{applyunitary}$ given in Fig. 7 is such that
@@
(a) The input to the algorithm is the description of a Gaussian state $|\psi\rangle$ which includes the overlap $r=\langle\alpha,\psi\rangle$ with a coherent state $|\alpha\rangle$, as well as a description $\Delta_{U}$ of a Gaussian unitary $U$.
(b) The reference state of the evolved state $U|\psi\rangle$ is $U|\alpha\rangle=|\alpha^{\prime}\rangle$. The inner product $r^{\prime}=\langle U\alpha,U\psi\rangle$ is equal to $r$ due to the unitarity of $U$. The algorithm returns $(\Gamma^{\prime},\alpha^{\prime},r)$, which is the description of the evolved state $U|\psi\rangle$.
Figure 8: An illustration of the algorithm $\mathsf{applyunitary}$ laid in Fig. 7 when $U$ is a displacement, a phase shift or a beamsplitter operator.
@@ 4.4 Computing descriptions of measured states
We give pseudocode for the algorithm $\mathsf{postmeasure}$ in Fig. 9, we prove that it computes the description of the post-measurement state in Theorem 4.8 and we represent the algorithm pictorially in Fig. 10.
@@
1:$(s_{A},s_{B})\leftarrow(\hat{d}(\alpha_{A}),\hat{d}(\alpha_{B}))$
2:$\Gamma^{\prime}\leftarrow$ evaluate Eq. 56
3:$s^{\prime}\leftarrow$ evaluate Eq. 57
4:$\alpha^{\prime}\leftarrow\hat{d}^{-1}(s^{\prime})$
5:$(\Gamma_{1},\Gamma_{2},\Gamma_{3})\leftarrow(\Gamma,\Gamma^{\prime},I)$triangleright covariance matrices of $\left(|\psi(\Delta)\rangle,|\psi^{\prime}\rangle,|\alpha^{\prime}\rangle\right)$
@@
7:$\lambda\leftarrow\alpha-\alpha^{\prime}$
8:$u\leftarrow\exp\left(i\operatorname{Im}\left(\alpha^{\prime T}\overline{\alpha}\right)\right)r$
9:$p\leftarrow\mathsf{prob}(\Delta,\beta)$
10:$v\leftarrow\pi^{k}\sqrt{p}$
11:$\overline{r^{\prime}}\leftarrow\mathsf{overlaptriple}(\Gamma_{1},\Gamma_{2},\Gamma_{3},d_{1},d_{2},d_{3},u,v,\lambda)$triangleright compute $\langle\alpha^{\prime},\psi^{\prime}\rangle$
@@
Figure 9: Given the description $\Delta\in\operatorname{\mathsf{Desc}}_{n}$ of a Gaussian state $|\psi\rangle\in\mathrm{Gauss}_{n}$ and $\beta\in\mathbb{C}^{k}$, the algorithm $\mathsf{postmeasure}$ returns the description of the post-measurement state $|\psi^{\prime}\rangle$ given in Eq. 129.
@@
The algorithm $\mathsf{postmeasure}:\operatorname{\mathsf{Desc}}_{n}\times\mathbb{C}^{k}$ given in Fig. 9 runs in time $O(n^{3})$. The algorithm is such that
@@
(a) The algorithm takes as input the description $\Delta$ of a Gaussian state $\psi=\psi(\Delta)$ that includes the overlap $r=\langle\alpha,\psi\rangle$ with the reference state $|\alpha\rangle$.
(b) The covariance matrix $\Gamma^{\prime}$ and displacement of the post-measurement state $|\psi^{\prime}\rangle$ are computed using the covariance matrix formalism. The reference state $|\alpha^{\prime}\rangle$ is the coherent state with the displacement of $|\psi^{\prime}\rangle$.
(c) The algorithm computes $\lambda\in\mathbb{C}^{n}$ and $\vartheta\in\mathbb{R}$ such that $D(\lambda)|\alpha\rangle=e^{i\vartheta}|\alpha^{\prime}\rangle$. Then, $\langle\alpha^{\prime},D(\lambda)\alpha\rangle=e^{i\vartheta}$ and $\langle\alpha^{\prime},D(\lambda)\psi\rangle=e^{i\vartheta}r$ are known.
(d) The overlap $\langle\psi,\psi^{\prime}\rangle=\|\psi^{\prime}\|=\pi^{k}\sqrt{p}$ with $p$ the measurement probability is known.
(e) Finally, the overlap $r^{\prime}=\langle\alpha^{\prime},\psi^{\prime}\rangle$ can be computed using the algorithm $\mathsf{overlaptriple}$, which corresponds to completing the triangle. The algorithm returns $(\Gamma^{\prime},\alpha^{\prime},r^{\prime})$.
Figure 10: Pictorial description of the algorithm $\mathsf{postmeasure}$.

### v2

@@ 3.2 Computing the overlap between Gaussian states
Figure 4: An illustration of the algorithm $\mathsf{overlap}$. a) The input to $\mathsf{overlap}$ consists of descriptions of two Gaussian states $\psi(\Delta_{1}),\psi(\Delta_{2})$ which includes $\alpha_{1},\alpha_{2}\in\mathbb{C}^{n}$ and the overlaps $r_{j}=\langle\alpha_{j},\psi(\Delta_{j})\rangle,j\in[2]$. b) The algorithm computes $\lambda\in\mathbb{C}^{n}$ and $\vartheta\in\mathbb{R}$ such that $D(\lambda)|\alpha_{1}\rangle=e^{i\vartheta}|\alpha_{2}\rangle$. This implies that $\langle\alpha_{2},D(\lambda)\alpha_{1}\rangle=e^{i\vartheta}$ and $\langle\psi(\Delta_{2}),D(\lambda)\alpha_{1}\rangle=e^{i\vartheta}\overline{r_{2}}$ are known. c) Lastly, the subroutine $\mathsf{overlaptriple}$ is applied to complete a triangle. This amounts to computing the inner product $w=\langle\psi(\Delta_{1}),\psi(\Delta_{2})\rangle$ which the algorithm returns.
@@ 3.3 Computing descriptions of evolved states
$\alpha^{\prime}_{k}$ $=\alpha_{k}\quad\text{for}\quad k\neq j\ .$ (105)
Because of this, we build an algorithmic subroutine called $\mathsf{squeezing}$ which, given the description $\Delta\in\operatorname{\mathsf{Desc}}_{n}$ of a Gaussian state $|\psi\rangle$, a parameter $z\in(0,+\infty]$ and $j\in[n]$, returns the description of the squeezed state $S_{j}(z)|\psi(\Delta)\rangle$. We call this subroutine in the algorithm $\mathsf{applyunitary}$ when $U$ is a single-mode squeezing operator.
First, we give the subroutine $\mathsf{squeezing}$, including pseudocode and associated claims in Lemma 3.6, as well as a pictorial description of the algorithm in Fig. 5.
@@
[…]
11:return $(\operatorname{\Gamma}^{\prime},\alpha^{\prime},r^{\prime})$
@@
The algorithm $\mathsf{squeezing}:\operatorname{\mathsf{Desc}}_{n}\times(0,+\infty)\times[n]\rightarrow\operatorname{\mathsf{Desc}}_{n}$ runs in time $O(n^{3})$. The algorithm is such that
$|\psi(\mathsf{squeezing}(\operatorname{\Delta},z,j))\rangle=S_{j}(z)|\psi(\operatorname{\Delta})\rangle$ (106)
for all $\operatorname{\Delta}\in\operatorname{\mathsf{Desc}}_{n},z\in\mathbb{R}$ and $j\in[n]$. That is, the output of the algorithm is a description of the state $S_{j}(z)|\psi(\Delta)\rangle$.
@@
Figure 5: An illustration of the algorithm $\mathsf{squeezing}$. Dashed lines correspond to inner products between the states at the vertices which are known to be non-zero but have not been computed. a) The input to the algorithm $\mathsf{squeezing}$ is the description of a Gaussian state $\psi$ which includes the overlap $r=\langle\alpha,\psi\rangle$, as well as $z\in\mathbb{R}$ and an index $j\in[n]$ associated to a single-mode squeezing operation $S_{j}(z)$. b) By unitarity of $S_{j}(z)$, the input data provides the inner product $\langle S_{j}(z)\alpha,S_{j}(z)\psi\rangle=\langle\alpha,\psi\rangle=r$. c) The reference state of the evolved state $S_{j}(z)|\psi\rangle$ is the coherent state $|\alpha^{\prime}\rangle$ with $\alpha^{\prime}\in\mathbb{C}^{n}$ such that $S(z)|\alpha\rangle=D(\alpha^{\prime})S(z)|0\rangle$. This is due to $S_{j}(z)|\psi\rangle$ having the same displacement vector $d(\alpha^{\prime})$ as $S_{j}(z)|\alpha\rangle$. d) The inner product $\langle\alpha^{\prime},S_{j}(z)\alpha\rangle$ is computed using Eq. 118. e) In the last step, the subroutine $\mathsf{overlaptriple}$ is used to compute $r^{\prime}=\langle\alpha^{\prime},S_{j}(z)\psi\rangle$. Thus, $(\operatorname{\Gamma}^{\prime},\alpha^{\prime},r^{\prime})$ is the description of $S_{j}(z)|\psi\rangle$.
@@
where $\alpha^{\prime}$ is the vector given in Eqs. 104 and 105 and initialized in Lines 4 and 5. The state $S_{j}(z)|\alpha\rangle$ has displacement vector $\hat{d}(\alpha^{\prime})$ and covariance matrix $\Gamma^{\prime\prime}=SS^{T}$. Since $|\alpha\rangle$ and $|\psi(\Delta)\rangle$ have the same displacement (by the definition of $\Delta$) and $S_{j}(z)$ maps states with the same displacement to states with the same displacement, then the state $S_{j}(z)|\psi(\Delta)\rangle$ has the same displacement $\hat{d}(\alpha^{\prime})$ as $|\alpha^{\prime}\rangle$. Then, $\Delta^{\prime}$ satisfies Item (i) in Definition 3.1.
It remains to show that the description $\Delta^{\prime}$ satisfies Item (ii) in Definition 3.1. This corresponds to showing that $r^{\prime}=\langle\alpha,S_{j}(z)\psi\rangle$. In Lines 6 and 7 we set
$d_{j}$ $=\hat{d}(\alpha^{\prime})\quad\text{for}\quad j\in[3]\ ,$ (111)
$(\Gamma_{1},\Gamma_{2},\Gamma_{3})$ $=(\Gamma^{\prime\prime},I,\Gamma^{\prime})\ ,$ (112)
@@
which we set as $r^{\prime}$ in Line 10. Hence, the description $\Delta^{\prime}$ satisfies Item (ii) in Definition 3.1.
@@
We give pseudocode for the algorithm $\mathsf{applyunitary}$ and in Theorem 3.7 we prove that it outputs the description of a Gaussian state evolved with a Gaussian unitary operation. We also give a diagram of the algorithm in Fig. 6 when the operation is a displacement, phase shit or beamsplitter. When the operation is a single-mode squeezing operator, we refer to Fig. 5 for a diagram of the algorithm.
@@
2:  $(z,j)\leftarrow\Delta_{U}$ $\triangleright$ $\Delta_{U}=(z,j)$ is a description of $U=S_{j}(z)$
@@
6:  $\beta\leftarrow\Delta_{U}$ $\triangleright$ $\Delta_{U}=\beta$ is a description of $U=D(\beta)$
@@
11:  $(\phi,j)\leftarrow\Delta_{U}$ $\triangleright$ $\Delta_{U}=(\phi,j)$ is a description of $U=F_{j}(\phi)$
@@
16:  $(\omega,j,k)\leftarrow\Delta_{U}$ $\triangleright$ $\Delta_{U}=(\omega,j,k)$ is a description of $U=B_{j,k}(\omega)$
@@
[…]
@@
The algorithm $\mathsf{applyunitary}$ is such that
@@
Figure 6: An illustration of the algorithm $\mathsf{applyunitary}$ when $U$ is a displacement, a phase shift or a beamsplitter operator. a) The input to the algorithm is the description of a Gaussian state $|\psi\rangle$ which includes the overlap $r=\langle\alpha,\psi\rangle$ with a coherent state $|\alpha\rangle$, as well as a description $\Delta_{U}$ of a Gaussian unitary $U$. b) The reference state of the evolved state $U|\psi\rangle$ is $U|\alpha\rangle=|\alpha^{\prime}\rangle$. The inner product $r^{\prime}=\langle U\alpha,U\psi\rangle$ is equal to $r$ due to the unitarity of $U$. The algorithm returns $(\Gamma^{\prime},\alpha^{\prime},r)$, which is the description of the evolved state $U|\psi\rangle$.
@@ 3.4 Computing descriptions of measured states
We give pseudocode for the algorithm $\mathsf{postmeasure}$, we prove that it computes the description of the post-measurement state in Theorem 3.8 and we represent the algorithm pictorially in Fig. 7.
@@
[…]
@@
The algorithm $\mathsf{postmeasure}:\operatorname{\mathsf{Desc}}_{n}\times\mathbb{C}^{k}$ runs in time $O(n^{3})$. The algorithm is such that
@@
Figure 7: Pictorial description of the algorithm $\mathsf{postmeasure}$. a) The algorithm takes as input the description $\Delta$ of a Gaussian state $\psi=\psi(\Delta)$ that includes the overlap $r=\langle\alpha,\psi\rangle$ with the reference state $|\alpha\rangle$. b) The covariance matrix $\Gamma^{\prime}$ and displacement of the post-measurement state $|\psi^{\prime}\rangle$ are computed using the covariance matrix formalism. The reference state $|\alpha^{\prime}\rangle$ is the coherent state with the displacement of $|\psi^{\prime}\rangle$. c) The algorithm computes $\lambda\in\mathbb{C}^{n}$ and $\vartheta\in\mathbb{R}$ such that $D(\lambda)|\alpha\rangle=e^{i\vartheta}|\alpha^{\prime}\rangle$. Then, $\langle\alpha^{\prime},D(\lambda)\alpha\rangle=e^{i\vartheta}$ and $\langle\alpha^{\prime},D(\lambda)\psi\rangle=e^{i\vartheta}r$ are known. d) The overlap $\langle\psi,\psi^{\prime}\rangle=\|\psi^{\prime}\|=\pi^{k}\sqrt{p}$ with $p$ the measurement probability is known. e) Finally, the overlap $r^{\prime}=\langle\alpha^{\prime},\psi^{\prime}\rangle$ can be computed using the algorithm $\mathsf{overlaptriple}$, which corresponds to completing the triangle. The algorithm returns $(\Gamma^{\prime},\alpha^{\prime},r^{\prime})$.